# RAS: Retrieval-And-Structuring for Knowledge-Intensive LLM Generation

### v2

@@ 4 Experiments
Test Datasets & Metrics & Compared Baselines. We conduct comprehensive evaluations on diverse knowledge-intensive tasks following previous studies [30, 31]. The evaluation encompasses three categories of datasets: (1) open-domain short-form generation datasets: TriviaQA [51], PopQA [52], and 2WikiMultihopQA [53]; (2) closed-set task datasets: PubHealth [54] and ARC-Challenge [55]; and (3) long-form generation datasets: ALCE-ASQA [56, 57] and ELI5 [58]. For evaluation metrics, we maintain consistency with prior work [30, 52, 31], employing “golden match” accuracy for PopQA and TriviaQA, token-level F1 score for 2WikiMultihopQA, accuracy for PubHealth and ARC-Challenge, and ROUGE-LSum alongside MAUVE score [59] for ASQA and ELI5. Our comparative analysis includes three baseline categories: models without retrieval augmentation, incorporating Claude 3.5 Sonnet as a state-of-the-art closed-source baseline; models with single retrieval over top-5 documents, including Claude 3.5 Sonnet, and SuRe [25], a leading retrieve-and-summarize method; and models with self-reflective retrieval, including leading approaches Self-RAG [30], RPG [31], and Search-R1 [60]. We place more details of metrics/datasets in Appendix D.
| Short-form | Closed-set | Long-form Generation | |||||||
| Model | TQA | 2WQA | PopQA | Pub | ARC | ASQA | ELI5 | ||
| (acc) | (F1) | (acc) | (acc) | (acc) | (rouge) | (mauve) | (rouge) | (mauve) | |
| Baselines w/o Retrieval | |||||||||
| Llama2 | 30.5 | 18.9 | 14.7 | 34.2 | 21.8 | 15.3 | 19.0 | 18.3 | 32.4 |
| Alpaca | 54.5 | – | 23.6 | 49.8 | 45.0 | 29.4 | 61.7 | – | – |
| Llama2 | 38.5 | 20.2 | 14.7 | 29.4 | 29.4 | 12.4 | 16.0 | 18.2 | 41.4 |
| Alpaca | 61.3 | – | 24.4 | 55.5 | 54.9 | 32.0 | 70.6 | – | – |
| Llama3 | 56.1 | 21.2 | 26.7 | 33.2 | 42.2 | 17.6 | 25.0 | 18.2 | 39.7 |
| ChatGPT | 74.3 | 24.8 | 29.3 | 70.1 | 75.3 | 36.2 | 68.8 | 22.8 | 32.6 |
| Sonnet-3.5 | 78.4 | 40.0 | 30.2 | 83.7 | 88.5 | 37.0 | 39.1 | 21.8 | 26.5 |
| Baselines w/ Retrieval (#docs=5 by default) | |||||||||
| Llama2 | 42.5 | 21.0 | 38.2 | 30.0 | 48.0 | 22.1 | 32.0 | 18.6 | 35.3 |
| Alpaca | 64.1 | – | 46.7 | 40.2 | 48.0 | 33.3 | 57.9 | – | – |
| Llama2 | 47.0 | 31.2 | 45.7 | 30.2 | 26.0 | 20.5 | 24.7 | 18.6 | 42.3 |
| Llama3 | 60.4 | 33.4 | 48.6 | 36.5 | 40.1 | 23.9 | 52.1 | 18.8 | 40.7 |
| Alpaca | 66.9 | – | 46.1 | 51.1 | 57.6 | 36.7 | 56.6 | – | – |
| RECOMP [61] | – | 32.4 | – | – | – | 36.5 | 76.0 | – | – |
| ChatGPT | 65.7 | 32.8 | 50.8 | 54.7 | 75.3 | 39.9 | 79.7 | 20.6 | 57.2 |
| Sonnet-3.5 | 69.1 | 41.9 | 51.5 | 49.1 | 88.6 | – | – | – | – |
| Sonnet-3.5 | 72.5 | 53.7 | 57.3 | 53.9 | 87.1 | 38.8 | 61.6 | 20.2 | 32.3 |
| SuRe [25] | 51.2 | 20.6 | 39.0 | 36.2 | 52.7 | 35.8 | 76.2 | 16.1 | 26.6 |
| SuRe | 72.3 | 38.1 | 53.6 | 57.2 | 79.6 | 36.0 | 74.2 | 19.2 | 51.6 |
| SuRe | 76.8 | 37.6 | 41.2 | 62.8 | 91.6 | 30.2 | 69.9 | 15.4 | 27.2 |
| Baselines w/ Self-Reflective Retrieval | |||||||||
| Self-RAG [30] | 66.4 | 25.1 | 54.9 | 72.4 | 67.3 | 35.7 | 74.3 | 17.9 | 35.6 |
| Self-RAG | 69.3 | – | 55.8 | 74.5 | 73.1 | 37.0 | 71.6 | – | – |
| RPG [31] | 65.1 | 33.6 | 56.0 | 73.4 | 65.4 | 37.6 | 84.4 | 19.1 | 46.4 |
| Search-R1 [60] | 63.8 | 35.2 | 45.7 | 69.4 | 64.4 | 35.2 | 72.3 | 18.4 | 37.1 |
| Retrieval-And-Structuring (RAS) | |||||||||
| RAS | 77.6 | 57.7 | 62.3 | 71.3 | 93.9 | 39.1 | 70.5 | 23.3 | 37.7 |
| RAS | 73.8 | 44.2 | 57.7 | 77.6 | 71.4 | 37.6 | 96.2 | 20.1 | 54.4 |
| RAS | 72.7 | 42.1 | 58.3 | 74.7 | 68.5 | 37.2 | 95.2 | 19.7 | 47.8 |
@@
Inference Setting. We evaluate RAS using both our trained open-source models (RAS, based on LLaMA-2-7B/LLaMA-3-8B and a Graph Transformer encoder [65]) and the closed-source Claude-3.5-Sonnet model under varied inference settings. For open-source models (RAS), we follow prior work [30, 31] and adopt zero-shot inference across all datasets. For the closed-source model (RAS), we apply few-shot inference with two exemplars for ASQA and ELI5, while using zero-shot inference for all other datasets. More details are provided in Appendix E.1 and E.2.
@@
During inference, if the initial planning output is not [NO_RETRIEVAL], we utilize the main query as the initial subquery . For PopQA and TriviaQA evaluation, we follow established settings [30, 66], incorporating top-five web search engine results as initial retrieved context . For ASQA and ELI5, we maintain methodological consistency with prior work [31, 30, 56], utilizing their predetermined five-document context. Under these conditions, we omit plan generation and text retrieval phases, implementing static inference [30] with five fixed iterations. For remaining datasets, we establish a maximum iteration count of five. Following previous studies, we employ Contriever-MS MARCO [67] as the primary dense retriever, with BM25 [68] serving as the retrieval mechanism for 2WikiMultihopQA. Across all retrieval processes, we maintain a consistent top- document selection of five. We present more details of inference settings in Appendix E.
4.1 Main Results
@@
Our performance evaluation, presented in Table 1, demonstrates that the LLaMA-2-7B/LLaMA-3-8B model fine-tuned with RAS outperforms existing open-source solutions, including Self-RAG [30] and RPG [31]. Notably, compared to the previous SOTA models, RAS shows a 9.7% improvement in short-form question-answering and a 7.9% gain in long-form generation tasks. Additionally, when applied to closed-source Sonnet-3.5 model, RAS consistently achieves superior results compared to single retrieval RAG approaches, including retrieve-and-summarize approach SuRe [25]. We find that sometimes (e.g., on TriviaQA and PubHealth) single-hop retrieval could not boost LLM’s performance, and even makes it worse, which demonstrates the necessity of on-demand retrieval, aligning with previous findings.
Our “role-swapping” study in Figure 3, where Sonnet-3.5 and RAS alternately perform the planning and answering tasks, reveals that performance is primarily bounded by the answering capability. For example, when Sonnet-3.5 handles planning and RAS handles answering, the system achieves 62.4% accuracy on the ARC-Challenge dataset. However, when Sonnet-3.5 is responsible for both planning and answering, accuracy increases to 93.9%. Notably, despite having 60 fewer parameters, our trained RAS achieves planning performance comparable to Sonnet-3.5, and even outperforms it in some cases.
Also, for long-form generation, our trained model achieved notably high MAUVE scores, which can be majorly attributed to our “text-to-triple” step, as demonstrated by our ablation study in Table 2.
4.2 Ablations & More Analysis
@@
| No GraphEncode | 70.2 | 38.4 | 66.4 | 33.1 | 85.0 |
| No LoRA | 71.5 | 37.8 | 54.8 | 32.8 | 84.8 |
| No Text-to-Triple | 70.4 | 38.2 | 71.4 | 36.2 | 73.8 |
| No Multi-Task | 68.6 | 39.2 | 65.5 | 36.7 | 88.9 |
@@
| No Retrieval | 56.9 | 27.4 | 69.0 | 31.3 | 70.6 |
| No GraphEncode | 68.8 | 38.7 | 67.3 | 36.5 | 93.6 |
| No Planning | 66.7 | 37.8 | 71.5 | 37.2 | 95.2 |
@@
Effect of LoRA and Multi-task Learning. Our experiments reveal that parameter-efficient training strategies significantly impact model performance. Using only graph token optimization without LoRA leads to substantial degradation (-11.8% on average). A similar observation can be made for “No Multi-Task” (e.g., 12.3% accuracy degradation on PubHealth), indicating the significance of jointly training the model on both action planning and answer generation tasks rather than optimizing for each task separately, supporting findings from prior work [31]. The complementary effects suggest that while graph-based knowledge representation is valuable, it needs to be combined with careful parameter tuning and multi-task learning to achieve optimal performance.
@@
Impact of Training Data Volume. Figure 4 demonstrates how training dataset size influences model performance across different tasks. Considering the computational efficiency, we sampled 2,000 instances each from TQA and 2WQA for evaluation, while maintaining the original sizes of other datasets. The results indicate that model performance generally improves with increased training data volume. Notably, our model achieves competitive performance even with limited training data - using only 5% (10K instances) of our full dataset, it surpasses previous state-of-the-art models on TQA, 2WQA, and ELI5. These results suggest both the robustness of our architectural design and the effectiveness of our data curation methodology. However, we observed an exception with the ELI5 dataset, where performance were inconsistent. This irregularity can be attributed to the inclusion of ASQA training data in our training set, following established setting from previous research [30, 31]. Among our test datasets, ASQA and ELI5 are unique in requiring long-form response generation. The periodic decline in ELI5 performance suggests that the model’s response generation began to align more closely with ASQA’s training data distribution, potentially at the expense of ELI5’s distinct characteristics.
| Triple Extractor | TQA | 2WQA | PopQA |
| Flan-T5-Large | 70.3 | 40.7 | 56.7 |
| LLaMA-3.2-3B | 72.7 | 42.1 | 58.3 |
| Claude-3.5-Sonnet | 73.8 | 44.5 | 60.1 |
Impact of Triple Extraction Quality We further examine how the quality of triple extraction influences the overall effectiveness of RAS. Specifically, we test three text-to-triples models: Flan-T5-Large, LLaMA-3.2-3B model and Claude-3.5-Sonnet where the first two are trained on WikiOfGraph data. As shown in Table 3, higher-quality triple extraction leads to consistent performance gains across all evaluated datasets. Notably, using Claude-3.5-Sonnet triples improves TriviaQA accuracy from 72.7% to 73.8%, boosts 2WikiMultihopQA F1 from 42.1 to 44.5, and raises PopQA accuracy from 58.3% to 60.1%. These results underscore that while RAS achieves strong performance even with lightweight extraction models, its effectiveness can be further enhanced by supplying more precise and complete structured knowledge during inference.
@@ 5 Conclusion
We presented RAS (Retrieval-And-Structuring), a novel framework that improves knowledge-intensive language generation by interleaving iterative retrieval with dynamic knowledge graph construction. RAS introduces key innovations, including structure-aware action planning and graph-augmented answering, enabling models to reason over progressively organized information tailored to each query. Extensive experiments across seven benchmarks demonstrate that RAS consistently outperforms existing methods, particularly in tasks requiring multi-step reasoning and factual consistency. Furthermore, the modular, graph-based architecture of RAS allows seamless integration with external knowledge sources and human feedback, making it well-suited for high-stakes domains where transparency and adaptability are essential. Limitations and directions for future work are discussed in Appendix A.

### v3

@@ 4.1 Settings
Test Datasets & Metrics & Compared Baselines. We conduct comprehensive evaluations on diverse knowledge-intensive tasks following previous studies (Asai et al., 2023; Lyu et al., 2024b). The evaluation encompasses three categories of datasets: (1) open-domain short-form generation datasets: TriviaQA (Joshi et al., 2017), PopQA (Mallen et al., 2022), and 2WikiMultihopQA (Ho et al., 2020); (2) closed-set task datasets: PubHealth (Zhang et al., 2023a) and ARC-Challenge (Clark et al., 2018); and (3) long-form generation datasets: ALCE-ASQA (Gao et al., 2023; Stelmakh et al., 2022) and ELI5 (Fan et al., 2019). For evaluation metrics, we maintain consistency with prior work (Asai et al., 2023; Mallen et al., 2022; Lyu et al., 2024b), employing “golden match” accuracy for PopQA and TriviaQA, token-level F1 score for 2WikiMultihopQA, accuracy for PubHealth and ARC-Challenge, and ROUGE-LSum alongside MAUVE score (Pillutla et al., 2021) for ASQA and ELI5. Our comparative analysis includes three baseline categories: models without retrieval augmentation, incorporating Claude 3.5 Sonnet as a state-of-the-art closed-source baseline; models with single retrieval over top-5 documents, including Claude 3.5 Sonnet, and SuRe (Kim et al., 2024b), a leading retrieve-and-summarize method; and models with self-reflective retrieval, including leading approaches Self-RAG (Asai et al., 2023), RPG (Lyu et al., 2024b) . We place more details of datasets and metrics in Appendices E.1 and E.2, respectively.
| Short-form | Closed-set | Long-form Generation | ||||||||
| Model/Method | TQA | 2WQA | PopQA | Pub | ARC | ASQA | ELI5 | |||
| (acc) | (F1) | (acc) | (acc) | (acc) | (rouge) | (mauve) | (rouge) | (mauve) | ||
| Closed-source | w/o Retrieval | |||||||||
| ChatGPT | 74.3 | 24.8 | 29.3 | 70.1 | 75.3 | 36.2 | 68.8 | 22.8 | 32.6 | |
| Sonnet-3.5 | 78.4 | 40.0 | 30.2 | 83.7 | 88.5 | 37.0 | 39.1 | 21.8 | 26.5 | |
| w/ Single Retrieval (#docs=5) | ||||||||||
| Sonnet-3.5 | 69.1 | 41.9 | 51.5 | 49.1 | 88.6 | n/a | n/a | n/a | n/a | |
| Sonnet-3.5 | 72.5 | 53.7 | 57.3 | 53.9 | 87.1 | 38.8 | 61.6 | 20.2 | 32.3 | |
| SuRe(Kim et al., 2024b) | 72.3 | 38.1 | 53.6 | 57.2 | 79.6 | 36.0 | 74.2 | 19.2 | 51.6 | |
| SuRe | 76.8 | 37.6 | 41.2 | 62.8 | 91.6 | 30.2 | 69.9 | 15.4 | 27.2 | |
| w/ Self-Reflective Retrieval | ||||||||||
| ReAct (Yao et al., 2023) | 73.4 | 53.7 | 55.0 | 62.2 | 89.2 | 38.8 | 61.6 | 20.2 | 32.3 | |
| IRCoT (Trivedi et al., 2023) | 74.7 | 54.9 | 53.2 | 59.4 | 92.0 | 38.8 | 61.6 | 20.2 | 32.3 | |
| Retrieval-And-Structuring (ours) | ||||||||||
| RAS | 77.6 | 57.7 | 62.3 | 71.3 | 93.9 | 39.1 | 70.5 | 23.3 | 37.7 | |
| Open-source | w/o Retrieval | |||||||||
| Llama2 | 30.5 | 18.9 | 14.7 | 34.2 | 21.8 | 15.3 | 19.0 | 18.3 | 32.4 | |
| Llama2 | 38.5 | 20.2 | 14.7 | 29.4 | 29.4 | 12.4 | 16.0 | 18.2 | 41.4 | |
| Llama3 | 56.1 | 21.2 | 26.7 | 33.2 | 42.2 | 17.6 | 25.0 | 18.2 | 39.7 | |
| w/ Single Retrieval (#docs=5) | ||||||||||
| Llama2 | 42.5 | 21.0 | 38.2 | 30.0 | 48.0 | 22.1 | 32.0 | 18.6 | 35.3 | |
| Llama2 | 47.0 | 31.2 | 45.7 | 30.2 | 26.0 | 20.5 | 24.7 | 18.6 | 42.3 | |
| Llama3 | 60.4 | 33.4 | 48.6 | 36.5 | 40.1 | 23.9 | 52.1 | 18.8 | 40.7 | |
| SuRe (Kim et al., 2024b) | 51.2 | 20.6 | 39.0 | 36.2 | 52.7 | 35.8 | 76.2 | 16.1 | 26.6 | |
| w/ Self-Reflective Retrieval | ||||||||||
| Self-RAG (Asai et al., 2023) | 66.4 | 25.1 | 54.9 | 72.4 | 67.3 | 35.7 | 74.3 | 17.9 | 35.6 | |
| Self-RAG | 69.3 | 26.9 | 55.8 | 74.5 | 73.1 | 37.0 | 71.6 | 18.7 | 38.5 | |
| RPG (Lyu et al., 2024b) | 65.1 | 33.6 | 56.0 | 73.4 | 65.4 | 37.6 | 84.4 | 19.1 | 46.4 | |
| ReAct (Yao et al., 2023) | 64.0 | 25.0 | 42.7 | 52.4 | 59.0 | 22.1 | 32.0 | 18.6 | 35.3 | |
| IRCoT (Trivedi et al., 2023) | 61.5 | 27.6 | 44.3 | 59.6 | 61.6 | 22.1 | 32.0 | 18.6 | 35.3 | |
| Retrieval-And-Structuring (ours) | ||||||||||
| RAS | 72.7 | 42.1 | 58.3 | 74.7 | 68.5 | 37.2 | 95.2 | 19.7 | 47.8 | |
| RAS | 73.8 | 44.2 | 57.7 | 77.6 | 71.4 | 37.6 | 96.2 | 20.1 | 54.4 | |
@@
Inference Setting. We evaluate RAS using both our trained open-source models (RAS, based on LLaMA-2-7B/LLaMA-3-8B and a Graph Transformer encoder (Shi et al., 2020)) and the closed-source Claude-3.5-Sonnet & Claude-4.5-Sonnet (Anthropic, 2025) under varied inference settings. For open-source models (RAS), we follow prior work (Asai et al., 2023; Lyu et al., 2024b) and adopt zero-shot inference across all datasets. For the closed-source model (RAS, RAS), we apply few-shot inference with two exemplars for ASQA and ELI5, while using zero-shot inference for all other datasets. More details are provided in Appendices F.1 and F.2.
@@
For PopQA and TriviaQA evaluation, we follow established settings (Asai et al., 2023; Luo et al., 2023a), incorporating top-five web search engine results as initial retrieved context . For ASQA and ELI5, we maintain methodological consistency with prior work (Lyu et al., 2024b; Asai et al., 2023; Gao et al., 2023), utilizing their predetermined five-document context. Under these conditions, we omit plan generation and text retrieval phases, implementing static inference (Asai et al., 2023) with five fixed iterations. For remaining datasets, we establish a maximum iteration count of five. Following previous studies, we employ Contriever-MS MARCO (Izacard et al., 2021) as the primary dense retriever, with BM25 (Robertson et al., 2009) serving as the retrieval mechanism for 2WikiMultihopQA. Across all retrieval processes, we maintain a consistent top- document selection of five. We present more details of inference settings in Appendix F.
4.2 Results
@@
Main Result. Our performance evaluation, presented in Table 1, demonstrates that the LLaMA-2-7B/LLaMA-3-8B model fine-tuned with RAS outperforms existing SFT-based open-source solutions, including Self-RAG (Asai et al., 2023) and RPG (Lyu et al., 2024b). Notably, compared to the earlier SOTA models, RAS shows a 9.7% improvement in short-form question-answering and a 7.9% gain in long-form generation tasks. Additionally, when applied to Claude-3.5-Sonnet, RAS consistently achieves superior results compared to single retrieval RAG approaches, including retrieve-and-summarize approach SuRe (Kim et al., 2024b). We find that sometimes (e.g., on TriviaQA and PubHealth) single-hop retrieval could not boost LLM’s performance, and even makes it worse, which demonstrates the necessity of on-demand retrieval, aligning with previous findings.
Although RAS integrates planning and answering in a unified framework, these components can be decoupled for greater flexibility. Our “role-swapping” study in Figure 4 demonstrates that performance is primarily constrained by answering capability rather than planning. When Sonnet-3.5 handles planning while RAS performs answering, the system achieves 62.4% accuracy on ARC-C, compared to 93.9% when Sonnet-3.5 performs both tasks. Despite having 60 fewer parameters, RAS achieves planning performance comparable to (and sometimes exceeding) Sonnet-3.5.
@@
| w/o GraphEncode | 70.2 | 38.4 | 66.4 | 33.1 | 85.0 |
| w/o LoRA | 71.5 | 37.8 | 54.8 | 32.8 | 84.8 |
| w/o Text-to-Triple | 70.4 | 38.2 | 71.4 | 36.2 | 73.8 |
| w/o Multi-Task | 68.6 | 39.2 | 65.5 | 36.7 | 88.9 |
@@
| w/o Retrieval | 56.9 | 27.4 | 69.0 | 31.3 | 70.6 |
| w/o GraphEncode | 68.8 | 38.7 | 67.3 | 36.5 | 93.6 |
| w/o Planning | 66.7 | 37.8 | 71.5 | 37.2 | 95.2 |
@@
Effect of LoRA and Multi-task Learning. Our experiments reveal that parameter-efficient training strategies significantly impact model performance. Using only graph token optimization without LoRA leads to substantial degradation (-11.8% on average). A similar observation can be made for “No Multi-Task” (e.g., 12.3% accuracy degradation on PubHealth), indicating the significance of jointly training the model on both action planning and answer generation tasks rather than optimizing for each task separately, supporting findings from prior work (Lyu et al., 2024b). The complementary effects suggest that while graph-based knowledge representation is valuable, it needs to be combined with careful parameter tuning and multi-task learning to achieve optimal performance.
@@
Impact of Training Data Volume. Figure 5 demonstrates how training dataset size influences model performance across different tasks. Considering the computational efficiency, we sampled 2,000 instances each from TQA and 2WQA for evaluation, while maintaining the original sizes of other datasets. The results indicate that model performance generally improves with increased training data volume. Notably, our model achieves competitive performance even with limited training data - using only 5% (10K instances) of our full dataset, it surpasses previous state-of-the-art models on TQA, 2WQA, and ELI5. These results suggest both the robustness of our architectural design and the effectiveness of our data curation methodology. However, we observed an exception with the ELI5 dataset, where performance were inconsistent. This irregularity can be attributed to the inclusion of ASQA training data in our training set, following established setting from previous research (Asai et al., 2023; Lyu et al., 2024b). Among our test datasets, ASQA and ELI5 are unique in requiring long-form response generation. The periodic decline in ELI5 performance suggests that the model’s response generation began to align more closely with ASQA’s training data distribution, potentially at the expense of ELI5’s distinct characteristics.
| Triple Extractor | Performance | Efficiency | ||
| TQA | 2WQA | PopQA | (tokens/s) | |
| Flan-T5-Large | 70.3 | 40.7 | 56.7 | 166.0 |
| LLaMA-3.2-3B | 72.7 | 42.1 | 58.3 | 4,885.3 |
| Claude-3.5-Sonnet | 73.8 | 44.5 | 60.1 | 68.2 |
Impact of Triple Extractor Selection We examine how triple extraction quality affects RAS performance using three models: Flan-T5-Large, LLaMA-3.2-3B, and Claude-3.5-Sonnet. As shown in Table 3, higher-quality extraction consistently improves results across datasets, with Claude-3.5-Sonnet achieving the best accuracy but at significantly lower efficiency (1.5e-2 sec/token vs 2.0e-4 for LLaMA-3.2-3B). These results demonstrate the accuracy-efficiency tradeoff in triple extraction, with LLaMA-3.2-3B providing the optimal balance for our experiments.
@@ 5 Conclusion
We presented RAS, a framework that dynamically constructs question-specific knowledge graphs through iterative retrieval and structured reasoning. Experiments across seven benchmarks show consistent improvements of up to 6.4% and 7.0% with open-source and proprietary LLMs, respectively. The modular architecture enables transparent reasoning and seamless integration with external knowledge sources. Limitations and future work are discussed in Appendix A. The statement of the LLM usage is placed in Appendix B. Broader Impacts, safeguards, and used assets are discussed in Appendix K. Our data and code can be found at: https://github.com/pat-jj/RAS.
Ethics Statement
This work does not involve human subjects, personal data, or sensitive user information. Our research focuses on algorithmic improvements for knowledge-intensive language model reasoning. We are mindful of potential ethical concerns, such as risks of misinformation or biased outputs when deploying retrieval-augmented LLMs. To mitigate these risks, we provide transparent descriptions of our methods, ablation studies on failure cases, and open-sourcing of code and models to facilitate community scrutiny. We also discuss broader impacts and safeguards in Appendix K.
Acknowledgment
Research was supported in part by the AI Institute for Molecular Discovery, Synthetic Strategy, and Manufacturing: Molecule Maker Lab Institute (MMLI), funded by U.S. National Science Foundation under Awards No. 2019897 and 2505932, NSF IIS 25-37827, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329. The research has used the Delta/DeltaAI advanced computing and data resource, supported in part by the University of Illinois Urbana-Champaign and through allocation #250851 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants OAC #2320345, #2138259, #2138286, #2138307, #2137603, and #2138296.